OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = -2x + 3?

pooja195 (pooja195):

Following points.....??? Where are they?

OpenStudy (anonymous):

(1, 1) (-1, 5) (-2, 3) (2, -1)

pooja195 (pooja195):

Since the directions say which is NOT a point on the graph we need to figure out which one of these do not lead to the right solution 1) y = -2x + 3 ; (1,1) 1=-2(1)+3 1=-2+3 1=1 Is this a solution? Yes or no?